A wet socks clinging to the inside of a washing machine drum which is spinning at a speed of




4.7 m/s. The radius of the drum is 30 cm. What is the minimum coefficient of static friction




between sock and the drum so that the sock will not slip down?

1
Expert's answer
2021-11-16T09:59:21-0500


The Newton’s second law equation

"F_N =ma_c \\

a_c = ω^2γ \\

f_s = mg \\

\mu_sF_N = mg \\

\mu_sm ω^2γ = mg \\

r= 0.3 \;m \\

ω = \frac{V}{r} = \frac{4.7}{0.3} = 15.667 \; rad/s \\

\mu_s = \frac{g}{ω^2γ} = \frac{9.8}{(15.667)^2 \times 0.3} = 0.133"

The minimum coefficient of static friction is 0.133